OpenStudy (anonymous):

I dont get it? whats the slope of this equation?: 8x-y=2

OpenStudy (anonymous):

u know y=mx+b?

OpenStudy (saifoo.khan):

eight.

OpenStudy (anonymous):

yeah, I just dont get how to set it up?

OpenStudy (saifoo.khan):

this is because the co-efficent of the x is 8

OpenStudy (anonymous):

solve for y pretty much. 8x-y=2 -y=-8x+2 y=8x-2 got it?

OpenStudy (anonymous):

yes!! thankyou :)